# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ 4 Proof of Theorem 2: eliminating the regular case
We have the following setting
  • $\mathfrak{g}$ is an absolutely simple Lie algebra
  • $\mathfrak{h},\mathfrak{l}$ are semisimple (and consist of simple ideals of non-compact type) subalgebras of $\mathfrak{g}$ s.t. the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ gives a standard compact Clifford-Klein form
  • The triple is not contained in Table 1 and thus there exists $0\neq X\in\mathfrak{m}_{0}$ orthogonal to $\mathfrak{h}\oplus\mathfrak{l}$.
@@
Denote by $\mathfrak{d}^{c}$ the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}^{c}.$ We have
$\mathfrak{d}^{c}=\mathfrak{m}_{0}^{c}\oplus(\mathfrak{a}_{h}^{\perp})^{c}\oplus(\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\oplus\mathfrak{a}_{h}^{c},$
where
$\mathfrak{m}_{0}^{c}=\mathfrak{t}^{c}\oplus\Sigma_{\alpha\in\Delta_{m}}\mathfrak{g}_{\alpha}.$
Notice that $\mathfrak{d}^{c}$ is a reductive subalgebra. Also if $\alpha\in\Delta_{0}^{\pm}$ and $\alpha|_{\mathfrak{a}_{h}}=0$ then (by Theorem 4) necessarily $\alpha(iX)=0.$ Consider the semisimple subalgebra
@@
The subalgebra $\mathfrak{d}^{c}_{s}$ is not simple or $\Phi:=\{\alpha\in\Delta_{0}^{\pm}\ |\ \alpha|_{\mathfrak{a}_{h}}=0\}=\emptyset.$
@@
Assume that $\Phi\neq\emptyset.$ If $\alpha\in\Delta_{m}$ is s.t. $\alpha(iX)\neq 0$ and $\beta\in\Phi$ then $\alpha\pm\beta$ is not a root since otherwise $\alpha\pm\beta\in\Phi$ and $(\alpha\pm\beta)(iX)\neq 0,$ which is impossible. Take any $\alpha_{1}\in\Phi.$ It follows that for any $\alpha\in\Delta_{m},$ $\alpha(iX)\neq 0$ we have $\alpha\perp\alpha_{1}.$ But now every root of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ is orthogonal to $iX$ or to $\alpha_{1}$ and so all roots are contained in two hyperplanes. Lemma 5 implies that $\mathfrak{d}_{s}^{c}$ is not simple. ∎
@@
We can decompose $\mathfrak{d}^{c}$ into a sum of three ideals $\mathfrak{d}^{c}=\mathfrak{d}_{1}\oplus\mathfrak{d}_{2}\oplus\mathfrak{a}_{h}^{c}$ such that $iX\in\mathfrak{d}_{1}$ and $\mathfrak{a}_{h}^{\perp},\ (\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\subset\mathfrak{d}_{2}.$
@@
First notice that $\mathfrak{a}_{h}^{c}$ is in the center of $\mathfrak{d}^{c}.$ If $\Phi=\emptyset$ then take $\mathfrak{d}_{1}:=\mathfrak{m}_{0}^{c},$ $\mathfrak{d}_{2}:=(\mathfrak{a}_{h}^{\perp})^{c}.$
@@
Assume that $\alpha_{1}\in\Phi\neq\emptyset.$ All roots of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ lie on two hypersurfaces in $(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c},$ $s_{iX}$ (orthogonal to $iX$) and $s_{\alpha_{1}}$ (orthogonal to $\alpha_{1}$). We see that if $\tilde{\mathfrak{d}}$ is a simple ideal of $\mathfrak{d}_{s}^{c}$ then (by Lemma 5) all roots of $(\tilde{\mathfrak{d}},\tilde{\mathfrak{d}}\cap(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c}))$ lie either in $s_{iX},$ or in $s_{\alpha_{1}}.$ Therefore we may take $\mathfrak{d}_{2}$ the be the sum of all simple ideals which roots lie in $s_{iX}$ plus $(\mathfrak{a}_{h}^{\perp})^{c}.$ Then $\mathfrak{d}_{1}$ is the ideal of $\mathfrak{m}_{0}^{c}$ complementary to $\mathfrak{d}_{2}\cap\mathfrak{m}_{0}^{c}.$
As $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$ then some Cartan subalgebra $\mathfrak{c}$ of $\mathfrak{g}^{c}$ normalizes $\mathfrak{h}^{c}.$ We may assume that $\mathfrak{a}_{h}^{c}\subset\mathfrak{c}.$ Indeed, notice that $\mathfrak{h}^{c}\cap\mathfrak{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}.$ Since any two Cartan subalgebras of $\mathfrak{h}^{c}$ are conjugate by an element of $H^{c}$ thus any Cartan subalgebra of $\mathfrak{h}^{c}$ is of the form $Ad_{h}(\mathfrak{c})\cap\mathfrak{h}^{c}$ for some $h\in H^{c}$ (where $H^{c}$ denotes a connected subgroup of $G^{c}$ corresponding to $\mathfrak{h}^{c}$). Obviously $Ad_{h}(\mathfrak{c})$ normalizes $\mathfrak{h}^{c}.$ As $\mathfrak{a}_{h}^{c}$ is in some Cartan subalgebra of $\mathfrak{h}^{c}$ therefore we may assume that
$\mathfrak{a}_{h}^{c}\subset\mathfrak{c}\subset\mathfrak{d}^{c}.$
Lemma 9.
We may assume that
$\sigma(\mathfrak{c})=\mathfrak{c},\ \ \theta(\mathfrak{c})=\mathfrak{c},\ \ \sigma(\mathfrak{d}_{2})=\mathfrak{d}_{2},\ \ \theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
Proof.
First notice that $\sigma(\mathfrak{a}_{h})=\mathfrak{a}_{h}$ and so $\sigma(\mathfrak{d}^{c})=\mathfrak{d}^{c}.$ Denote by $\mathfrak{d}^{\sigma}$ the real form of $\mathfrak{d}^{c}$ given by $\sigma.$ We see that $\mathfrak{d}^{\sigma}$ is the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}.$ Put $\mathfrak{c}_{h}:=\mathfrak{c}\cap\mathfrak{h}^{c}.$ After conjugating $\mathfrak{c}$ by an element of the centralizer of $\mathfrak{a}_{h}$ in $H^{c}$ we have $\sigma(\mathfrak{c}_{h})=\mathfrak{c}_{h}.$ Therefore $\mathfrak{c}_{h}^{\sigma}$ (the real form of $\mathfrak{c}_{h}$ w.r.t. $\sigma$) is in $\mathfrak{d}^{\sigma}.$ We take a $\sigma$-stable Cartan subalgebra $\mathfrak{c}^{\prime}$ of the normalizer of $\mathfrak{h}^{c}$ in $\mathfrak{d}^{c}$ containing $\mathfrak{c}_{h}^{\sigma}.$ We have $[\mathfrak{c}^{\prime},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}$ and $\sigma(\mathfrak{c}^{\prime})\subset\mathfrak{c}^{\prime},$ as desired. By the similar argument (by conjugating $(\mathfrak{c}^{\prime})^{\sigma}$ inside $G$) we have $\theta(\mathfrak{c}^{\prime})=\mathfrak{c}^{\prime}.$
Notice that as $\sigma(X)=\theta(X)=X$ we have $\sigma(\mathfrak{d}_{2})=\theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
Lemma 10.
We have $\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}=\{0\}$ and so $\Phi\neq\emptyset.$
Proof.
Let $Y\in\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}.$ For every $1\leq i\leq k$ we have
$[Y,S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
$[Y,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i}))]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
where $\mathfrak{n}_{h}^{\perp}$ is the orthogonal complement of $\mathfrak{n}_{h}$ in $\mathfrak{n}$ w.r.t. $B_{\theta}(\bullet,\bullet)=B(\bullet,\theta(\bullet)).$ It follows that $Y$ centralizes all roots of $\Delta_{p}^{\pm}$ and $\Delta_{n}^{\pm}.$ By Lemma 7 every root of $\Delta_{0}^{\pm}$ is of the form $\alpha_{1}+\alpha_{2}$ for some $\alpha_{1},\alpha_{2}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ But now $Y$ centralizes $\mathfrak{n}$ and so $Y=0.$
Notice that
$\mathfrak{d}^{\sigma}\subset\mathfrak{k}_{d}\oplus\mathfrak{m}_{0}\oplus\mathfrak{p}_{d}\oplus\mathfrak{a}_{h},$
$\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
$\mathfrak{p}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
where
$\mathfrak{p}_{d}=\mathfrak{a}_{h}^{\perp}\oplus\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})-\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))-\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\},$
$\mathfrak{k}_{d}=\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))+\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\}.$
Previous Lemmas and the fact that $\textrm{rank}(\mathfrak{g})>\textrm{rank}(\mathfrak{m}_{0}\oplus\mathfrak{a}_{h})$ imply the following corollary.
@@
Obviously $\mathfrak{a}_{h}\subset\mathfrak{c}^{\sigma}.$ Take any $Y\in\mathfrak{p}_{d}\cap\mathfrak{c}_{\sigma}.$ Then $Y$ is conjugate by an element $k^{-1}$ of $K_{d}$ (connected closed subgroup of $K$ corresponding to reductive subalgebra $\mathfrak{k}_{d}$) to $\mathfrak{a}_{h}^{\perp}.$ As $\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$ thus it centralizes $iX$ and $\mathfrak{a}_{h}$ and so for any $g\in K_{d}$ $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$). Also $\mathfrak{a}_{h}^{\perp}$ is orthogonal to $\textrm{Ad}_{g}\mathfrak{h}.$ We can repeat the argument of the proof of Lemma 10 to $\textrm{Ad}_{k}Y$ and $\textrm{Ad}_{k}\mathfrak{h}.$
Because $\theta(\mathfrak{d}_{2}^{\sigma})=\mathfrak{d}_{2}^{\sigma}$ therefore $\mathfrak{d}_{2}^{\sigma}\subset\mathfrak{k}.$ As every element of a Lie algebra of compact type is in some Cartan subalgebra we may conjugate $\mathfrak{c}^{\sigma}$ by an element of $\mathfrak{d}_{2}^{\sigma}$ so that
$0\neq\tilde{Y}:=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}))\in\mathfrak{c}^{\sigma},$
where $\alpha\in\Phi.$ As $\mathfrak{d}_{2}^{\sigma}$ centralizes $iX$ and $\mathfrak{a}_{h}$ thus for any $g\in D_{2}^{\sigma}$ (connected closed subgroup of $G$ corresponding to reductive subalgebra $\mathfrak{d}_{2}^{\sigma}$) $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$). Thus Theorem 4 holds for $\textrm{Ad}_{g}\mathfrak{h}.$

### v3

@@ 4 Proof of Theorem 2: eliminating the regular case
4.1 Assumptions
In this section we consider the following triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$:
  1. 1.
    $\mathfrak{g}$ is an absolutely simple Lie algebra,
  2. 2.
    $\mathfrak{h},\mathfrak{l}$ are absolutely simple (of non-compact type) subalgebras of $\mathfrak{g}$ and the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ gives a standard compact Clifford-Klein form,
  3. 3.
    the triple is not contained in Table 1 and thus there exists $0\neq X\in\mathfrak{m}_{0}$ orthogonal to $\mathfrak{h}+\mathfrak{l}$,
  4. 4.
@@ 4.1 Assumptions
In this section we prove the "regular" part of Theorem 2, that is, the following.
Theorem 7.
Assume that neither $G/H$ nor $H$ is compact. Under the assumptions of this section the pair $(\mathfrak{g},\mathfrak{h})$ cannot admit compact standard Clifford-Klein forms.
4.2 Plan of the proof
The plan of the proof is as follows.
  1. 1.
    The assumptions that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form and that there exists $0\neq X\not\in\mathfrak{h}+\mathfrak{l}$ yield the conclusions of Theorem 4, that is, some particular properties of the spaces $Z_{h}$ and $Z_{l}$ with bases $S_{h}^{i}$ and $S_{l}^{i}$ (statement 1), the relation between $S_{h}^{i}$ and $\Sigma_{h}$ ( statement 2), and the form of the base of the complexified real root spaces $\mathfrak{h}_{\gamma}^{c}$ (statement 3).
  2. 2.
    The assumption of the regularity of $\mathfrak{h}^{c}$ together with the assumption $X\not\in\mathfrak{h}+\mathfrak{l}$ yields the existence of a Cartan subalgebra $\mathfrak{c}\subset\mathfrak{g}^{c}$ such that $\mathfrak{a}_{h}\subset\mathfrak{c}$ and an element $0\not=Y\in\mathfrak{c}$ such that $Y\not\in\mathfrak{m}_{0}$, but $Y\in\mathfrak{z}_{\mathfrak{g}}(\mathfrak{h})$ and $[Y,iX]=0$.
  3. 3.
    Statements 1) and 2) are contradictory, and, therefore, $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$, as required.
4.3 Deriving the contradiction between 1) and 2)
Let us show first, how to derive the required contradiction. Put
$Z_{p}=\sum_{\alpha\in\Delta_{p}^{\pm}}\mathfrak{g}_{\alpha},Z_{n}=\sum_{\alpha\in\Delta_{n}^{\pm}}\mathfrak{g}_{\alpha},O=\sum_{\alpha\Delta_{0}}\mathfrak{g}_{\alpha}.$
By 2),
$[Y,iX]=0,[Y,\mathfrak{a}_{h}]=0.$ (3)
This implies
$[Y,Z_{n}]\subset Z_{n},[Y,Z_{p}]\subset Z_{p},[Y,O]\not\subset(Z_{p}+Z_{n})\setminus\{0\}.$
Indeed, $\mathfrak{g}_{\alpha}\subset Z_{p}$ if and only if $\alpha(iX)>0$ and $\alpha(A^{+})>0$ for any $A^{+}$ from the interior of the positive Weyl chamber in $\mathfrak{a}_{h}$ (we consider Weyl chambers in $\mathfrak{a}_{h}$). Let $U\in\mathfrak{g}_{\alpha}\subset Z_{p}$. The Jacobi identity and (3) imply
$[iX,[Y,U]]=\alpha(iX)[Y,U],\,\text{and}\,\,[iX,[Y,A^{+}]]=\alpha(A^{+})[Y,U].$
Thus, $[Y,U]$ is contained in a sum of root spaces determined by roots which are positive on $iX$ and on $A^{+}$, as required. Similar arguments work for $Z_{n}$ and $O$.
Since $[Y,\mathfrak{h}^{c}]=0$, by Theorem 4 (statements 2 and 3), $[Y,S_{h}^{i}+Q_{i}]=0$ for all $i=1,...,k$. Theorem 4 (statement 1) implies
$[Y,x_{\alpha_{i}}]=0,\,\,\text{for all}\,\,\alpha_{i}\in\Delta_{p}^{\pm}.$
It follows that $[Y,\sigma(x_{\alpha_{i}})]=0$ as well, and, therefore, $[Y,x_{\alpha_{j}}]=0$ for any $\alpha_{j}\in\Delta^{\pm}_{n}$. Lemma 4 yields $[Y,\mathfrak{a}]=0$ which contradicts 2), that is, the assumption $Y\not\in\mathfrak{m}_{0}$.
4.4 Constructing $Y\in\mathfrak{c},Y\not\in\mathfrak{m}_{0},Y\in\mathfrak{z}_{\mathfrak{g}}(\mathfrak{h}),[Y,iX]=0$
We will construct the required $Y$ in the centralizer $\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})$.
4.4.1 The structure of $\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})$
One easily sees that
$\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})=\mathfrak{d}^{c}=\mathfrak{m}_{0}+(\mathfrak{a}_{h}^{\perp})^{c}+\mathfrak{a}_{h}+\sum_{\alpha\in\Delta_{0},\alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha},$
where $(\mathfrak{a}_{h}^{\perp})^{c}$ is the complexification of the orthogonal complement to $\mathfrak{a}_{h}$ in $\mathfrak{a}$. This is a reductive subalgebra in $\mathfrak{g}^{c}$. Since $\mathfrak{h}^{c}$ is regular, there exists a Cartan subalgebra $\mathfrak{c}\subset\mathfrak{g}^{c}$ such that $[\mathfrak{c},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}.$ Clearly, $\mathfrak{h}^{c}\cap\mathfrak{c}$ is a Cartan subalgebra in $\mathfrak{h}^{c}$. Since any two Cartan subalgebras in $\mathfrak{h}^{c}$ are conjugate by elements $\operatorname{Ad}(h)$ (where $h$ is an element in the connected subgroup $H^{c}\subset G^{c}$ corresponding to $\mathfrak{h}^{c}$), any Cartan subalgebra of $\mathfrak{h}^{c}$ is of the form $\operatorname{Ad}(h)(\mathfrak{c})\cap\mathfrak{h}^{c}$. It easily follows that without loss of generality one may assume that $\mathfrak{a}_{h}^{c}\subset\mathfrak{c}$. Since $\mathfrak{d}^{c}$ centralizes $\mathfrak{a}_{h}$ one finally obtains
$\mathfrak{a}_{h}^{c}\subset\mathfrak{c}\subset\mathfrak{d}^{c}.$ (4)
Lemma 5.
The Cartan subalgebra $\mathfrak{c}$ can be chosen in a way to satisfy $(4)$ and to be $\sigma$- and $\theta$-invariant.
Proof.
Clearly, $\sigma(\mathfrak{a}_{h})=\mathfrak{a}_{h}$. It follows that $\sigma(\mathfrak{d}^{c})=\mathfrak{d}^{c}.$ Denote by $\mathfrak{d}^{\sigma}$ the real form of $\mathfrak{d}^{c}$ given by $\sigma.$ We see that $\mathfrak{d}^{\sigma}$ is the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}.$ Put $\mathfrak{c}_{h}:=\mathfrak{c}\cap\mathfrak{h}^{c}.$ After conjugating $\mathfrak{c}$ by an element of the centralizer of $\mathfrak{a}_{h}$ in $H^{c}$ we get $\sigma(\mathfrak{c}_{h})=\mathfrak{c}_{h}.$ Therefore $\mathfrak{c}_{h}^{\sigma}$ (the real form of $\mathfrak{c}_{h}$ with respect to $\sigma$) is in $\mathfrak{d}^{\sigma}.$ Take a $\sigma$-stable Cartan subalgebra $\mathfrak{c}^{\prime}$ of the normalizer of $\mathfrak{h}^{c}$ in $\mathfrak{d}^{c}$ containing $\mathfrak{c}_{h}^{\sigma}.$ One checks that $[\mathfrak{c}^{\prime},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}$ and $\sigma(\mathfrak{c}^{\prime})\subset\mathfrak{c}^{\prime},$ as desired. By a similar argument (conjugating $(\mathfrak{c}^{\prime})^{\sigma}$ inside $G$) we get $\theta(\mathfrak{c}^{\prime})=\mathfrak{c}^{\prime}.$
Lemma 6.
The following equality is satisfied:
$\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}=\{0\}.$
Proof.
Let $Y\in\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}.$ For every $1\leq i\leq k$ we get the following:
$[Y,S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})]=0,$
$[Y,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i}))]=0,$
This easily follows from the fact that for any $i=1,...,k$, vectors
$S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})\,\,\text{and}\,\,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i})$
belong to $\mathfrak{h}$, and the properties of the Killing form. Indeed, one can notice that $[Y,h]\perp\mathfrak{h}$ for any $h\in\mathfrak{h}$. On the other hand, $[Y,h]\in\mathfrak{h}$ since $Y$ normalizes $\mathfrak{h}$. Now observe that since $i$ takes all values from $1$ to $k$, $Y$ normalizes all root vectors of $\Delta_{p}^{\pm}$ and $\Delta_{n}^{\pm}.$ By Lemma 4 every root of $\Delta_{0}^{\pm}$ is of the form $\alpha_{1}+\alpha_{2}$ for some $\alpha_{1}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and $\alpha_{2}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\cup\Delta_{m}.$ This implies that $Y$ centralizes $\mathfrak{n}$ and so $Y=0.$
Consider the semisimple part of the reductive subalgebra $\mathfrak{d}^{c}$, that is
@@ 4.4.1 The structure of $\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})$
At least one of the following is valid:
  1. 1.
    $\mathfrak{d}^{c}_{s}$ is not simple,
  2. 2.
    $iX$ is in the center of $\mathfrak{d}^{c}.$
@@
If $iX$ is in the center of $\mathfrak{d}^{c}$ then the proof is complete. So assume that $\alpha(iX)\neq 0$ for some $\alpha\in\Delta_{m}.$ By Lemma 6, $\Phi\not=\emptyset$. If $\beta\in\Phi$ then $\alpha\pm\beta$ is not a root since otherwise $\alpha\pm\beta\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and $(\alpha\pm\beta)|_{\mathfrak{a}_{h}}=0,$ which is impossible (by Theorem 4). Take any $\alpha_{1}\in\Phi.$ It follows that for any $\gamma\in\Delta_{m},$ $\gamma(iX)\neq 0$ we have $\gamma\perp\alpha_{1}.$ But now every root of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ is orthogonal to $iX$ or to $\alpha_{1}$ and so all roots are contained in two hyperplanes (but not all roots are in the first one, nor in the second one). Lemma 3 implies that $\mathfrak{d}_{s}^{c}$ is not simple. ∎
@@
Lie algebra $\mathfrak{d}^{c}$ decomposes into a direct sum of Lie algebras
$\mathfrak{d}^{c}=\mathfrak{d}_{1}\oplus\mathfrak{d}_{2}\oplus\mathfrak{a}_{h}^{c}$
such that
  1. 1.
    $iX\in\mathfrak{d}_{1}$
  2. 2.
    $\mathfrak{a}_{h}^{\perp}\oplus\Sigma_{\alpha\in\Delta_{0},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha}\subset\mathfrak{d}_{2}.$
  3. 3.
    $\sigma(\mathfrak{d}_{2})\subset\mathfrak{d}_{2},\theta(\mathfrak{d}_{2})\subset\mathfrak{d}_{2}$.
@@
Note that $\mathfrak{a}_{h}^{c}$ is in the center of $\mathfrak{d}^{c}.$ If $iX\in\mathfrak{z}(\mathfrak{d}^{c})$, then put $\mathfrak{d}_{1}:=\textrm{Span}_{\mathbb{C}}(iX)$ and take the ideal $\mathfrak{d}_{2}$ complementary to $\textrm{Span}_{\mathbb{C}}(iX)\oplus\mathfrak{a}_{h}^{c}.$ Otherwise, assume that $\alpha_{1}\in\Phi\neq\emptyset$ and $\alpha(iX)\neq 0$ for some $\alpha\in\Delta_{m}.$ Define hyperplanes $s_{iX}$ and $s_{\alpha_{1}}$ in $(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c},$ orthogonal to $iX$ and $\alpha_{1},$ respectively. All roots of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ lie in these two hyperplanes. We see that if $\tilde{\mathfrak{d}}$ is a simple ideal of $\mathfrak{d}_{s}^{c}$, then (by Lemma 3) all roots of $(\tilde{\mathfrak{d}},\tilde{\mathfrak{d}}\cap(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c}))$ lie either in $s_{iX},$ or in $s_{\alpha_{1}}.$ Therefore we may take $\mathfrak{d}_{2}$ to be the sum of all simple ideals whose roots lie in $s_{iX}+(\mathfrak{a}_{h}^{\perp})^{c}.$ Then $\mathfrak{d}_{1}$ is the ideal of $\mathfrak{m}_{0}^{c}$ complementary to $\mathfrak{d}_{2}\cap\mathfrak{m}_{0}^{c}.$
Finally, the third part of the Corollary follows, since $\sigma(X)=X$ and $\theta(X)=X$. ∎
Let $\mathfrak{k}_{d}$ denote the subalgebra of $\mathfrak{g}$ generated by vectors
$x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))+\theta(i(x_{\alpha}-\sigma(x_{\alpha}))),\forall\alpha\in\Phi,$
and $\mathfrak{p}_{d}$ denote the subspace in $\mathfrak{g}$ generated by $\mathfrak{a}_{h}^{\perp}$ and all vectors of the form
$x_{\alpha}+\sigma(x_{\alpha})-\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))-\theta(i(x_{\alpha}-\sigma(x_{\alpha}))),\forall\alpha\in\Phi.$
From the very definition of $\mathfrak{k}_{d}$ and $\mathfrak{p}_{d}$, as well as the definition of $\mathfrak{d}^{c}$ and taking into consideration Corollary 2 one obtains the inclusions
$\mathfrak{k}_{d}\oplus\mathfrak{p}_{d}\subset\mathfrak{d}_{2}^{\sigma}\subset\mathfrak{d}^{\sigma}\subset\mathfrak{k}_{d}\oplus\mathfrak{p}_{d}\oplus\mathfrak{m}_{0}\oplus\mathfrak{a}_{h}.$ (5)
@@
It is straightforward to see that $\mathfrak{a}_{h}\subset\mathfrak{c}^{\sigma}.$ Take any $Y\in\mathfrak{p}_{d}\cap\mathfrak{c}^{\sigma}.$ Let $K_{d}$ be a connected closed (reductive) subgroup of $K$ corresponding to $\mathfrak{k}_{d}$. Since $\mathfrak{a}_{h}^{\perp}$ is maximal abelian in $\mathfrak{p}_{d}$, there is an element $k\in K_{d}$ such that $\operatorname{Ad}_{k}(Y)\in\mathfrak{a}^{\perp}_{h}$.
Since $\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$, it centralizes $iX$ and $\mathfrak{a}_{h}$ and so for any $g\in K_{d}$ the subalgebra $\operatorname{Ad}_{g}\mathfrak{h}$ is well embedded (since $\operatorname{Ad}_{g}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\operatorname{Ad}_{g}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\operatorname{Ad}_{g}\mathfrak{h})=\operatorname{Ad}_{g}\mathfrak{h}$). Also $\mathfrak{a}_{h}^{\perp}$ is orthogonal to $\operatorname{Ad}_{g}\mathfrak{h}.$ Now, one can repeat the argument of the proof of Lemma 6 (taking the appropriate Lie brackets) applying it to $\operatorname{Ad}_{k}Y$ and $\operatorname{Ad}_{k}\mathfrak{h}.$
4.4.2 Construction of $Y$
Let $D_{2}^{\sigma}$ denote the connected closed subgroup of $G$ corresponding to $\mathfrak{d}_{2}^{\sigma}$.